# Rotation of the Globular Cluster Population of the Dark Matter Deficient Galaxy NGC 1052-DF4: Implication for the total mass

### v1

@@ 1 Introduction
Ultra-diffuse galaxies (UDGs) are defined by their extremely low matter densities and luminosities, with a surface brightness from 25 to 28 mag arcsec-2 and a relatively large scale radius of about $>1.5$ kpc (2015ApJ...798L..45V; 2015ApJ...807L...2K). According to 2016ApJ...828L...6V, some UDGs appear to have very high dark matter fractions with very little stellar material. Alternatively, other UDGs appear to possess much less dark matter than expected (e.g. 2018Natur.555..629V).
@@
The potential lack of dark matter in NGC 1052-DF2 and NGC 1052-DF4 has prompted investigation from several angles. Some have focused on determining the distance to these galaxies (2019MNRAS.486.1192T; 2020ApJ...895L...4D; 2021ApJ...909..179S; 2021MNRAS.504.1668Z), with the possibility that a smaller distance would bring the properties of these galaxies inline with the overall galaxy populations. Others have considered the estimation of their intrinsic velocity dispersion (2018ApJ...864L..18V; 2018ApJ...859L...5M; 2019A&A...625A..77F; 2019ApJ...874L...5V), as well as the question of whether the associated globular cluster population rotates (2020MNRAS.491L...1L).
@@
However, it is important to note that the claim of dark matter deficiency is based on dynamical modeling that assumes no rotation (2020MNRAS.491L...1L). Such rotation shifts the balance between pressure and rotational support, and hence, neglecting it will bias the inferred dark matter mass (2018ApJ...863L..15W; 2022EPJC...82..935L).Inspired by the work of 2020MNRAS.491L...1L on NGC 1052-DF2, a similar statistical analysis of NGC 1052-DF4’s globular clusters was undertaken. The primary objective of this study is to determine whether the globular cluster population rotates.
@@ 2 Data
Previous authors used the Hubble Space Telescope to discover the compact objects associated with NGC 1052-DF4 (van Dokkum et al. 2019). Using the technique described in 2018ApJ...856L..30V, SExtractor (1996AAS..117..393B) was used to assess the overall magnitudes, colors, and Full Width Half Maximum (FWHM) sizes of compact objects in the images. 2018ApJ...856L..30V stated that compact objects with an $I_{814}$ magnitude smaller than 23, a $V_{606}-I_{814}$ colour within the range between 0.20 and 0.43, and a FWHM value between 0.12” and 0.30” are considered to be likely globular clusters. Hence, by applying this method, 2019ApJ...874L...5V found that, out of 11 candidates, only seven meet these criteria, with a mean overall magnitude with a $I_{814}$ value of 22.10 and an rms spread of 0.39 mag. These seven objects were subsequently confirmed as being associated with NGC 1052-DF4 using spectra from the Low Resolution Imaging Spectrograph on the Keck I telescope. Information about the seven confirmed globular clusters of NGC 1052-DF4 was available from 2019ApJ...874L...5V. For our purposes, the data consists of the positions of the clusters on the sky, the measured velocities of the clusters along the line of sight, and the uncertainties on these line-of-sight velocities. The positions and velocities of these seven clusters are shown in Figure 1.
@@ 3 Models and Bayesian Inference
The kinematic model used to analyze the NGC1052-DF2 galaxy (2020MNRAS.491L...1L) is still used in this study. The line-of-sight velocity due to rotation is given by:
@@
$A$ Rotational Amplitude Uniform(0, 20) km s-1
@@
[…]
@@ 4.3 Model 3: Rotational Model with Alternative Priors
$A$ Rotational Amplitude $(10^{1-|\textnormal{Student}-t(0,0,2)|})\sigma$ km s-1
@@ 5 Mass of the Galaxy
The posterior distribution of NGC 1052-DF4’s estimated mass with three different rotational inclinations is shown in Figure 5. This figure suggests that the estimated mass of NGC 1052-DF4 has right-skewed posterior distributions for all three inclination angles, with the majority of the probability falling between 0 and $10^{8}$ solar masses. Moreover, the estimated mass of the NGC 1052-DF4 can reach as high as 6×$\times 10^{8}$ solar masses. There is no substantial difference between the three density curves with different inclinations. This is reasonable as the amplitude, $A$, was very small or possibly zero, which caused the velocity dispersion, $\sigma$, to dominate in the equation 8 and result in the inclination becoming irrelevant.
@@ 6 Discussion
Based on equation 1, the amplitude is directly proportional to the rotational velocity. Also, from equation 5, it can be seen that the rotational velocity and the velocity dispersion, $\sigma$, are directly related to the mass of the galaxy. Consequently, a low amplitude value (around 0) and a slightly larger value of $\sigma$ indicate that the globular cluster rotational velocities will likewise be small, resulting in a very low estimated mass of the NGC 1052-DF4 galaxy, which is about the same as in the previous study (2019ApJ...874L...5V). Hence, these results further imply that the NGC 1052-DF4 galaxy contains little dark matter, which is consistent with the result of a newly published paper from 2023ApJ...957....6S. As mentioned in Section 1, 2024arXiv240204304G found that NGC 0152-DF4 exhibits tidal tails, which explains the low quantity of dark matter in the galaxy.
@@ 7 Conclusions
In this paper, we tested the existence of a rotating element by analysing the rotational features of the seven globular clusters in NGC 1052-DF4. After comparing the marginal likelihood and calculating the Bayes factor of the Non-Rotational Model versus the Rotational Model, we find that the Non-Rotational Model is weakly preferred by a Bayes Factor of about 2.5. The outcome of the Rotational Model with Alternative Priors provides more evidence that the globular cluster populations of NGC 1052-DF4 do not rotate significantly. The result from this study is consistent with the previous study (2019ApJ...874L...5V). Furthermore, this result supports the conclusion that there is little dark matter in the NGC 1052-DF4 galaxy since mass estimates to not need to be modified to take rotation into account. This result is also consistent with the recent study from 2024arXiv240204304G and reinforces the conclusion from 2023ApJ...957....6S.

### v2

@@ 1 Introduction
Ultra-diffuse galaxies (UDGs) are defined by their extremely low matter densities and luminosities, with a surface brightness from 25 to 28 mag arcsec-2 and a relatively large scale radius of about >1.5>1.5 kpc 2015ApJ...798L..45V; 2015ApJ...807L...2K. According to 2016ApJ...828L...6V, some UDGs appear to have very high dark matter fractions with very little stellar material. Alternatively, other UDGs appear to possess much less dark matter than expected (2018Natur.555..629V, e.g.).
@@
The potential lack of dark matter in NGC 1052-DF2 and NGC 1052-DF4 has prompted investigation from several angles. Some have focused on determining the distance to these galaxies 2019MNRAS.486.1192T; 2020ApJ...895L...4D; 2021ApJ...909..179S; 2021MNRAS.504.1668Z, with the possibility that a smaller distance would bring the properties of these galaxies inline with the overall galaxy populations. Others have considered the estimation of their intrinsic velocity dispersion 2018ApJ...864L..18V; 2018ApJ...859L...5M; 2019A&A...625A..77F; 2019ApJ...874L...5V, as well as the question of whether the associated globular cluster population rotates 2020MNRAS.491L...1L.
@@
However, it is important to note that the claim of dark matter deficiency is based on dynamical modeling that assumes no rotation 2020MNRAS.491L...1L. Such rotation shifts the balance between pressure and rotational support, and hence, neglecting it will bias the inferred dark matter mass 2018ApJ...863L..15W; 2022EPJC...82..935L.Inspired by the work of 2020MNRAS.491L...1L on NGC 1052-DF2, a similar statistical analysis of NGC 1052-DF4’s globular clusters was undertaken. The primary objective of this study is to determine whether the globular cluster population rotates.
@@ 2 Data
Previous authors used the Hubble Space Telescope to discover the compact objects associated with NGC 1052-DF4 (van Dokkum et al. 2019). Using the technique described in 2018ApJ...856L..30V, SExtractor 1996AAS..117..393B was used to assess the overall magnitudes, colors, and Full Width Half Maximum (FWHM) sizes of compact objects in the images. 2018ApJ...856L..30V stated that compact objects with an $I_{814}$ magnitude smaller than 23, a $V_{606}-I_{814}$ colour within the range between 0.20 and 0.43, and a FWHM value between 0.12” and 0.30” are considered to be likely globular clusters. Hence, by applying this method, 2019ApJ...874L...5V found that, out of 11 candidates, only seven meet these criteria, with a mean overall magnitude with a $I_{814}$ value of 22.10 and an rms spread of 0.39 mag. These seven objects were subsequently confirmed as being associated with NGC 1052-DF4 using spectra from the Low Resolution Imaging Spectrograph on the Keck I telescope. Information about the seven confirmed globular clusters of NGC 1052-DF4 was available from 2019ApJ...874L...5V. For our purposes, the data consists of the positions of the clusters on the sky, the measured velocities of the clusters along the line of sight, and the uncertainties on these line-of-sight velocities. The positions and velocities of these seven clusters are shown in Figure 1.
@@ 3 Models and Bayesian Inference
The kinematic model used to analyze the NGC1052-DF2 galaxy 2020MNRAS.491L...1L is still used in this study. The line-of-sight velocity due to rotation is given by:
@@
AA Rotational Amplitude Uniform(0, 20) km s-1
@@
$\mathcal{Z}=\int\mathcal{L}(\omega)\times\pi(\omega)\,d\omega,$ (5)
@@ 4.3 Model 3: Rotational Model with Alternative Priors
AA Rotational Amplitude $(10^{1-|\textnormal{Student}-t(0,0,2)|})\sigma$ km s-1
@@ 5 Mass of the Galaxy
[…]
@@ 6 Discussion
Based on equation 1, the amplitude is directly proportional to the rotational velocity. Also, from equation 5, it can be seen that the rotational velocity and the velocity dispersion, $\sigma$, are directly related to the mass of the galaxy. Consequently, a low amplitude value (around 0) and a slightly larger value of $\sigma$ indicate that the globular cluster rotational velocities will likewise be small, resulting in a very low estimated mass of the NGC 1052-DF4 galaxy, which is about the same as in the previous study 2019ApJ...874L...5V. Hence, these results further imply that the NGC 1052-DF4 galaxy contains little dark matter, which is consistent with the result of a newly published paper from 2023ApJ...957....6S. As mentioned in Section 1, 2024arXiv240204304G found that NGC 0152-DF4 exhibits tidal tails, which explains the low quantity of dark matter in the galaxy.
@@ 7 Conclusions
In this paper, we tested the existence of a rotating element by analysing the rotational features of the seven globular clusters in NGC 1052-DF4. After comparing the marginal likelihood and calculating the Bayes factor of the Non-Rotational Model versus the Rotational Model, we find that the Non-Rotational Model is weakly preferred by a Bayes Factor of about 2.5. The outcome of the Rotational Model with Alternative Priors provides more evidence that the globular cluster populations of NGC 1052-DF4 do not rotate significantly. The result from this study is consistent with the previous study 2019ApJ...874L...5V. Furthermore, this result supports the conclusion that there is little dark matter in the NGC 1052-DF4 galaxy since mass estimates to not need to be modified to take rotation into account. This result is also consistent with the recent study from 2024arXiv240204304G and reinforces the conclusion from 2023ApJ...957....6S.